# Complementary Time–Space Tradeoff for Self-Stabilizing Leader Election: Polynomial States Meet Sublinear Time

### v1

@@ 3.6 Finding a Target
We design the sub-protocol FindTarget by slightly modifying the collision-detection protocol CollisionDetectionWithBounds (abbr. CDWB) of Araya and Sudo [8]. Protocol CDWB detects rank collisions in $O(n^{3/2}\cdot\sqrt{\log n})$ expected interactions using $O(n\cdot\mathrm{poly}(\log n))$ states, and never reports a false positive when starting from a correctly initialized configuration. We modify CDWB so that, if a duplicate rank (i.e., a rank shared by two or more agents) exists, all agents agree on one such rank and store it in $\mathtt{target}$ within $O(n^{3/2}\log n)$ interactions with probability $1-o(1)$. If no duplicate exists, FindTarget may set $\mathtt{target}$ arbitrarily (we default to 1).
@@
Protocol CDWB relies on two assumptions, both of which we satisfy here. First, it requires rough knowledge of the population size $n$ (asymptotically tight lower and upper bounds); we instead provide exact knowledge of $n$. Second, it assumes a unique leader; we maintain a unique leader $a_{L}$ via the LS-LE protocol $\mathcal{P}_{{\textsc{LSLE}}}$ (line 1) (see Section 3.2).
@@
Hence, we obtain FindTarget as follows. Upon switching to mode $\mathsf{F}$, each agent initializes all CDWB variables and sets $\mathtt{target}=1$. Thereafter, on each interaction they execute CDWB. If a collision is detected, they compute a duplicate rank $r$ and overwrite $\mathtt{target}$ with $r$. In phase $T_{2}-2$—the final phase for this sub-protocol—agents stop running CDWB and instead execute an epidemic on $\mathtt{target}$, so that the maximum duplicate rank found so far propagates to the entire population.
@@
We now lower-bound the success probability of FindTarget by $1-o(1)$. Araya and Sudo [8] proved that for any integer $k\in[1,\Theta(\log n)]$, CDWB detects a collision (if exists) within $O(k\cdot n^{3/2}\cdot\sqrt{\log n})$ interactions with probability $1-(2/3)^{k}$. (See the proof of Lemma 5 in [8].) Since agents remain in phases $[T_{1},T_{2}-3]$ for $\Theta(n^{3/2}\cdot\log n)$ interactions, choosing $k=\sqrt{\log n}$ gives success probability $1-o(1)$.
@@
Starting from any configuration $C\in\mathcal{C}_{\mathrm{init}}(\mathsf{F})$ with at least one duplicate rank, the population under $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ reaches a configuration where every agent has $\mathtt{target}=r$ within $O(n^{3/2}\log n)$ interactions with probability $1-o(1)$.
3.7 Proof of Main Theorem
@@
We first bound the number of states. In $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$, the largest domain arises from the variable $\mathtt{list}$ in ${\textsc{Detect}}(r,\rho)$, which ranges over all subsets of $R_{\mathrm{name}}=[1,\rho^{2}]\setminus\{r\}$ of size at most $\rho$. Hence the size of the doamin is at most
$\sum_{i=0}^{\rho}\binom{\rho^{2}}{i}\leq\sum_{i=0}^{\rho}(\rho^{2})^{i}=O(\rho^{2\rho})=O(2^{2\rho\lg\rho}).$
All other variables range over polynomially sized domain. Thus, the total number of states is at most $O(2^{2\rho\lg\rho})\cdot\operatorname{poly}(n)=2^{2\rho\lg\rho+O(\log n)}$.
Next we bound the stabilization time. By Lemma 6, it suffices to show that, from any initial configuration, the population enters $\mathcal{S}$ within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions with probability $\Omega(1)$; the expected-time bound then follows.
By Lemma 4, within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions w.h.p. the system reaches some $C_{0}\in\mathcal{I}$. If $C_{0}$ has no duplicate rank, then $\mathtt{det}=0$ and $\mathtt{mode}=\bot$ in $C_{0}$, so $C_{0}\in\mathcal{S}$ and we are done. Otherwise, suppose $C_{0}$ contains at least one duplicate rank.
By Lemma 3, within $O(n\log n)$ interactions w.h.p. the execution enters $\mathcal{C}_{\mathrm{init}}(\mathsf{F})$ from $C_{0}\in\mathcal{I}$. Then by Lemma 8, within $O(n^{3/2}\log n)$ interactions with probability $1-o(1)$, FindTarget causes all agents to agree on a duplicate rank $r$, reaching $\mathcal{C}_{\mathrm{init}}(\mathsf{D},r)$. From there, Lemma 5 implies that within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions with probability $\Omega(1)$ every agent raises $\mathtt{det}=1$ and the execution enters $\mathcal{C}_{\mathrm{init}}(\mathsf{R})$. Finally, by Lemma 7, within $O(n^{3/2}\log n)$ interactions w.h.p. the population assigns all $n$ agents distinct ranks, and by Lemma 3 another $O(n^{3/2}\log n)$ interactions suffice to reach phase $T_{4}$. Then, all agents reset $\mathtt{mode}$ and $\mathtt{det}$ by epidemic for $\mathtt{reset}$, thus entering $\mathcal{S}$.
By a union bound over these phases, the probability of entering $\mathcal{S}$ within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions is still $\Omega(1)$. This completes the proof. ∎
@@ 4 Conclusion
In this paper, we have presented a new time–space tradeoff for self-stabilizing leader election (SS-LE) in the population protocol model under exact knowledge of the population size $n$. Independently of the recent protocol of Austin et al. [9], we showed:
  • For any integer $1\leq\rho\leq\sqrt{n}$, there is an SS-LE protocol that stabilizes in $O(\nicefrac{{n}}{{\rho}}\cdot\log\rho)$ expected time using $2^{2\rho\lg\rho+O(\log n)}$ states.
  • By choosing $\rho=\Theta\left(\nicefrac{{\log n}}{{\log\log n}}\right)$, this yields the first polynomial-state protocol with sublinear expected time.
Compared to the protocol of Austin et al., ours uses drastically fewer states whenever the target expected time is at least $\Theta(\sqrt{n}\log n)$.
Our result complements prior work by filling the gap between the linear-state, linear-time regime and the super-exponential-state, optimal-time regime. A remaining limitation is that the restriction $\rho\leq\sqrt{n}$ prevents us from achieving $o(\sqrt{n}\log n)$ expected time; by contrast, the protocol of Austin et al. can reach the optimal $\Theta(\log n)$ time at the cost of super-exponential state size.
@@
  • Does there exist an SS-LE protocol using only polynomially many states that stabilizes in $O(n^{1-\epsilon})$ time for some constant $\epsilon>0$? Our time–space tradeoff achieves only $O\left(n/\operatorname{poly}(\log n)\right)$ time under polynomial-state constraints.
@@ Acknowledgments
Appendix A Omitted Proofs
@@
In ${\textsc{AssignRanks}}_{\rho}$, the $n$ agents are eventually splited to $\rho$ deputies and $n-\rho>n/2$ recipients. Before assigning the $n$ agents unique ranks, each recipient must meet at least one deputy, which requires at least $\sum_{i=1}^{n-\rho}\binom{n}{2}/(2i\rho)=\Theta(\nicefrac{{n^{2}}}{{r}}\cdot\log n)$ expected interactions, i.e., $\Theta(\nicefrac{{n}}{{r}}\cdot\log n)$ time. ∎
@@
Sudo et al. [28] prove that for any constant $\tau\geq 1$, there exists a set $\mathcal{C}_{\mathrm{safe}}$ of configurations of $\mathcal{P}_{{\textsc{LSLE}}}(\tau)$ such that (i) starting from an arbitrary configuration, the population under $\mathcal{P}_{{\textsc{LSLE}}}(\tau)$ reaches a configuration in $\mathcal{C}_{\mathrm{safe}}$ within $O(\tau n\log n)$ expected interactions, and (ii) starting from any configuration in $\mathcal{C}_{\mathrm{safe}}$, the population goes back to a safe configuration after $\Theta(\tau n\log n)$ interactions and retains a unique leader during the period with probability $1-n^{-\tau}$ (See the proof of Lemma 15 in Appendix A.1 of [28] for details). The lemma immediately follows from this fact and the union bound, adjusting $\tau=\Theta(1)$ sufficiently. ∎

### v2

@@ 8 Finding a Target
In this section, we design the subprotocol FindTarget by slightly modifying the collision-detection protocol CollisionDetectionWithBounds of [AS25], abbreviated CDWB. Protocol CDWB detects rank collisions in $O(n^{3/2}\cdot\sqrt{\log n})$ expected interactions using $O(n\cdot\mathrm{poly}(\log n))$ states, and never reports a false positive when starting from a correctly initialized configuration. We modify CDWB so that, if a duplicate rank (i.e., a rank shared by two or more agents) exists, all agents agree on one such rank and store it in $\mathtt{target}$ within $O(n^{3/2}\log n)$ interactions with probability 1o(1)1-o(1). If no duplicate exists, FindTarget may set $\mathtt{target}$ arbitrarily (we default to 1).
@@
Protocol CDWB relies on two assumptions, both of which are satisfied in our setting. First, it requires approximate knowledge of the population size $n$ (asymptotically tight lower and upper bounds), whereas we provide exact knowledge of $n$. Second, it assumes a unique leader to operate a single-leader phase clock [AAE08]; we maintain such a leader $a_{L}$ via the LS-LE protocol $\mathcal{P}_{{\textsc{LSLE}}}$ (line 8) and run the phase clock (line 17) (see Section 5), allowing the computed phase to be reused by CDWB.
@@
Hence, we obtain FindTarget as follows. Upon switching to mode $\mathsf{F}$, each agent sets $\mathtt{target}=1$ (and resets auxiliary variable $\mathtt{parity}$ to $T_{1}\bmod{2}$.) Thereafter, on each interaction they execute CDWB. If a collision is detected, they compute a duplicate rank $r$ and overwrite $\mathtt{target}$ with $r$. In phase $T_{2}-2$, i.e., the final phase for this sub-protocol, agents stop running CDWB and instead execute an epidemic on $\mathtt{target}$, so that the maximum duplicate rank found so far propagates to the entire population. For completeness, we include the pseudocode of FindTarget in the appendix (Algorithm 4).
@@
We now lower‐bound the success probability of FindTarget by 1o(1)1-o(1). [AS25] proved that for any integer $k\in[1,\Theta(\log n)]$, CDWB detects a collision (if exists) within $O(k\cdot n^{3/2}\cdot\sqrt{\log n})$ interactions with probability $1-(2/3)^{k}$. (See the proof of Lemma 5 in [AS25].) Since agents remain in phases $T_{1},T_{2}-3]$ for $\Theta(n^{3/2}\cdot\log n)$ interactions, choosing $k=\sqrt{\log n}$ gives success probability 1o(1)1-o(1).
@@
For any configuration $C$, let $R(C)$ denote the set of duplicate ranks appearing in $C$. Starting from any $C\in\mathcal{C}_{\mathrm{init}}(\mathsf{F})$ with $|R(C)|\geq 1$, the population under $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$ reaches, within $O(n^{3/2}\log n)$ interactions and with probability 1o(1)1-o(1), a configuration in which every agent stores the same value $r\in R(C)$ in the variable $\mathtt{target}$.
9 Proof of Main Theorem
@@
We first bound the number of states. In $\mathcal{P}_{{\textsc{SSRK}}}(\rho)$, the largest domain arises from the variable $\mathtt{list}$ in ${\textsc{Detect}}(r,\rho)$, which ranges over all subsets of $R_{\mathrm{name}}=[1,\rho^{2}]\setminus\{r\}$ of size at most $\lfloor\rho\lg\rho\rfloor$. Hence the size of the domain is at most
$\sum_{i=0}^{\lfloor\rho\lg\rho\rfloor}\binom{\rho^{2}}{i}\leq\sum_{i=0}^{\lfloor\rho\lg\rho\rfloor}(\rho^{2})^{i}=O(\rho^{2\lfloor\rho\lg\rho\rfloor})=O(2^{2\rho\lg^{2}\rho}).$
All other variables range over polynomially sized domain. Thus, the total number of states is at most $O(2^{2\rho\lg^{2}\rho})\cdot\operatorname{poly}(n)=2^{2\rho\lg^{2}\rho+O(\log n)}$.
Next, we bound the stabilization time. By Lemma 5, for any $T=\operatorname{poly}(n)$, starting from any configuration the population reaches, within $O(n\log n)$ expected interactions, a configuration $C$ from which a unique leader is preserved for at least $T$ interactions w.h.p. Therefore, by Lemma 4, it suffices to show that, starting from any such configuration $C$, the population reaches a safe configuration within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions with probability $\Omega(1)$ (see Definition 4 for the definition of safe configurations); the expected-time bound then follows. (Note that Lemma 4 does not require the unique-leader assumption.)
By Lemma 7, within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions w.h.p. the system reaches some $C_{0}\in\mathcal{I}$. If $C_{0}$ has no duplicate rank, then $\mathtt{det}=0$ and $\mathtt{mode}=\bot$ in $C_{0}$, so $C_{0}$ is safe and we are done. Otherwise, suppose $C_{0}$ contains at least one duplicate rank.
By Lemma 6, within $O(n\log n)$ interactions w.h.p. the execution enters $\mathcal{C}_{\mathrm{init}}(\mathsf{F})$ from $C_{0}\in\mathcal{I}$. Then by Lemma 9, within $O(n^{3/2}\log n)$ interactions with probability 1o(1)1-o(1), FindTarget causes all agents to agree on a duplicate rank $r$, reaching $\mathcal{C}_{\mathrm{init}}(\mathsf{D},r)$. From there, Lemma 3 implies that within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions with probability $\Omega(1)$ every agent raises $\mathtt{det}=1$ and the execution enters $\mathcal{C}_{\mathrm{init}}(\mathsf{R})$. Finally, by Lemma 8, within $O(n^{3/2}\log n)$ interactions w.h.p. the population assigns all $n$ agents distinct ranks, and by Lemma 6 another $O(n^{3/2}\log n)$ interactions suffice to reach phase $T_{4}$. Then, all agents reset $\mathtt{mode}$ and $\mathtt{det}$ by epidemic for $\mathtt{reset}$, thus reaching a safe configuration.
By a union bound over these phases, the probability of reaching a safe configuration within $O(\nicefrac{{n^{2}}}{{\rho}}\cdot\log\rho)$ interactions is still $\Omega(1)$. This completes the proof. ∎
@@ 10 Conclusion
In this paper, we have presented a new time–space tradeoff for self-stabilizing leader election (SS-LE) in the population protocol model under exact knowledge of the population size $n$. Independently of the recent protocol of [ABF+25], we showed:
  • For any integer $2\leq\rho\leq\sqrt{n}$, there is an SS-LE protocol that stabilizes in $O(\nicefrac{{n}}{{\rho}}\cdot\log\rho)$ expected time using $2^{2\rho\lg^{2}\rho+O(\log n)}$ states.
  • By choosing $\rho=\Theta\left(\nicefrac{{\log n}}{{\log^{2}\log n}}\right)$, this yields the first polynomial-state protocol with sublinear expected time.
Compared to the protocol of [ABF+25], ours uses drastically fewer states whenever the target expected time is at least $\Theta(\sqrt{n}\log n)$.
Our result complements prior work by filling the gap between the linear-state, linear-time regime and the super-exponential-state, optimal-time regime. A remaining limitation is that the restriction $\rho\leq\sqrt{n}$ prevents us from achieving $o(\sqrt{n}\log n)$ expected time; by contrast, the protocol of [ABF+25] can reach the optimal $\Theta(\log n)$ time at the cost of super-exponential state size.
@@
  • Does there exist an SS-LE protocol using only polynomially many states that stabilizes in $O(n^{1-\epsilon})$ time for some constant $\epsilon>0$? Our time–space tradeoff achieves only $O\left(n/\operatorname{poly}(\log n)\right)$ time under polynomial‐state constraints.
@@ Acknowledgments
Appendix A Omitted Proof and Pseudocode
@@
In ${\textsc{AssignRanks}}_{\rho}$, the $n$ agents are eventually split into $\rho$ deputies and $n-\rho>n/2$ recipients. Before assigning unique ranks to all $n$ agents, each recipient must meet at least one deputy. Therefore, the expected number of interactions to assign distinct ranks to all agents is at least
$\sum_{i=1}^{n-\rho}\frac{\binom{n}{2}}{i\,\rho}=\Omega\left(\tfrac{n^{2}\log n}{\rho}\right),$
i.e., $\Omega(\nicefrac{{n}}{{\rho}}\cdot\log n)$ expected time. ∎
@@
[SEIM21] prove that for any constant $\tau\geq 1$, there exists a set $\mathcal{C}_{\mathrm{safe}}$ of configurations of $\mathcal{P}_{{\textsc{LSLE}}}(\tau)$ such that (i) starting from an arbitrary configuration, the population under $\mathcal{P}_{{\textsc{LSLE}}}(\tau)$ reaches a configuration in $\mathcal{C}_{\mathrm{safe}}$ within $O(\tau n\log n)$ expected interactions, and (ii) starting from any configuration in $\mathcal{C}_{\mathrm{safe}}$, the population goes back to a safe configuration after $\Theta(\tau n\log n)$ interactions and retains a unique leader during the period with probability $1-n^{-\tau}$ (See the proof of Lemma 15 in Appendix A.1 of [SEIM21] for details). The lemma immediately follows from this fact and the union bound, adjusting $\tau=\Theta(1)$ sufficiently. ∎